OpenStudy (anonymous):

A crate of mass 50kg is dragged along a horzontal floor by a constant force of magnitude 400N acting at an angle of thita upwards from the horizontal total resistance to motion of the crate has constant magnitude 250N.The crate starts from rest at the point O and passes point P with a speed of 2m/s.The distance OP is 20m.For the crate`s motion from O to P. Given that the increase in kinetic energy is 100J.And Work done against resistance to motion of the crate is 5000joules A)Find the value of thita

OpenStudy (saifoo.khan):

im back.

OpenStudy (saifoo.khan):

a) \[\frac12 m v^2\]

OpenStudy (saifoo.khan):

a)\[\frac12 (50)(2)^2\]

OpenStudy (saifoo.khan):

b) WD = f * d = 250 * 20

OpenStudy (anonymous):

YES

OpenStudy (saifoo.khan):

c)